Could you help me? I need to find all asymptotes. F (x)=(x^2+3x-10)/(x^2+8x+15) thanks.
Answer by CubeyThePenguin(3113)

f(x) = (x^2 + 3x - 10)/(x^2 + 8x + 15)
= ((x + 5)(x - 2))/((x + 3)(x + 5))
= (x - 2)/(x + 3)

vertical asymptote: denominator can't be 0
x = -3

horizontal asymptote: compare the degrees of the first term in numerator and denominator
y = 1/1 = 1